OpenStudy (anonymous):

I need help with this question it is of permutation How many three digits number can be formed with each digit being distinct help ?

OpenStudy (anonymous):

@shamim The question is how many 3 digit numbers can be formed in which there is no repetition of any number that is each digit is different ? for example 678 how many such numbers can be made with 10 natural numbers

OpenStudy (uri):

@ipm1988 Do you have answer choices?

OpenStudy (anonymous):

no @uri

OpenStudy (uri):

9P3=9*8*7 Btw @shamim How do we get 9? :)

OpenStudy (anonymous):

@shamim would it not be 7*9*10 @uri

OpenStudy (uri):

@shamim Waat.

OpenStudy (uri):

@ash2326

OpenStudy (ash2326):

There are ten digits 0-9 Now hundredth place can't be 0, it'll make the no. a 2 digit So we have 9 choices for the the 100th place. Do you get this @ipm1988

OpenStudy (anonymous):

@uri _ _ _ these three blanks represent the numbers so if we take natural number then units can be filled in 10 ways that is 0 to 9 , tens can be filled in 9 ways and hundreds can be filled in 7 ways so it should be 7*9*10

OpenStudy (anonymous):

@ash2326 @uri @shamim

OpenStudy (anonymous):

@ash2326 but when you generally fill up numbers you start from right that is from units placce

OpenStudy (ash2326):

it doesn't matter, we can work both ways

OpenStudy (anonymous):

so answer should be 630 @ash2326

OpenStudy (ash2326):

Do you have the answer given?

OpenStudy (uri):

@ipm1988 You're in which grade?

OpenStudy (anonymous):

yes the answer 648 = 9*9*8 that is what is confsuing me @uri @shamim @ash2326

OpenStudy (ash2326):

If we work from the units place, first place can be filled in 10 ways, second can be in 9. But how do you have 7 for the hundredth? It's not that simple you have to consider cases where 0 is not being used, and where you have 0 being used.

OpenStudy (anonymous):

@uri I am in first year college just brushing up on some concepts

OpenStudy (ash2326):

@ipm1988 Are you ok with the last post I made?

OpenStudy (anonymous):

@ash please explain

OpenStudy (anonymous):

@ash2326 please explain I mean I get this you can either use whole or natural numbers zero i also understand but explain

OpenStudy (ash2326):

ok. Suppose we don't use 0, we have 9 no.s at hand _ _ _ unit's place- 9 ten's- 8 hundred's- 7 Total distinct no.s - 9 x 8 x 7=504

OpenStudy (uri):

That means i was right..and @shamim too

OpenStudy (uri):

9P3=9*8*7=504

OpenStudy (ash2326):

Uri we have to also consider the cases where 0 is being used, so the answer would be larger than this

OpenStudy (ash2326):

@ipm1988 do you get my last post?

OpenStudy (shamim):

10p3

OpenStudy (anonymous):

@ash2326 yes thanks